OpenStudy (anonymous):

An equation and the steps to solve it are shown below: 3x - one over two (8x - 2) = 4 Step 1: 3x - 4x + 1 = 4 Step 2: x + 1 = 4 Step 3: x + 1 - 1 = 4 - 1 Step 4: x = 3 Which step was the error made and why?

OpenStudy (debbieg):

Is this your step 1?\[\frac{ 3x-1 }{2(8x-2) }=4\]I'm having trouble deciphering... an image or some more parantheses would help.

OpenStudy (cwrw238):

yes - its pretty vague as it stands

OpenStudy (anonymous):

3x - one over two (8x - 2) = 4 Step 1: 3x - 4x + 1 = 4 Step 2: x + 1 = 4 Step 3: x + 1 - 1 = 4 - 1 Step 4: x = 3

OpenStudy (anonymous):

it's wanting to know the error

OpenStudy (debbieg):

There's no need to spell out the numbers, lol. OK, you're saying "3x - one over two (8x - 2) = 4" but that isn't at all what you have written as: Step 1: 3x - 4x + 1 = 4 so just going with the enumerated steps you show: Step 1: 3x - 4x + 1 = 4 Step 2: x + 1 = 4 3x-4x is not =x... 3x-4x=-x Step 3: x + 1 - 1 = 4 - 1 Step 4: x = 3

OpenStudy (anonymous):

I copied and paste it it did that by it's self.

OpenStudy (anonymous):

so it would be step 1?

OpenStudy (debbieg):

Well, the error is the result given in step 2 of the operation stated in step 1. Whichever step you want to call it, that's the error.

OpenStudy (anonymous):

thank you

OpenStudy (debbieg):

Wait.... i think I see what you mean,t he starting equation is: \[3x-\frac{ 1 }{ 2 }(8x-2)=4\] Then you get to step 1. It doesn't change my explanation, I just understand where the equation started now, lol.

OpenStudy (debbieg):

you're welcome.